# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v1

@@ Abstract.
The healthcare landscape is evolving, with patients seeking more reliable information about their health conditions, treatment options, and potential risks. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust doctors and hospital staff, highlighting the need for expert-endorsed health information. However, the pressure on experts has led to reduced communication time, impacting information sharing. To address this gap, we propose CataractBot, an experts-in-the-loop chatbot powered by large language models (LLMs). Developed in collaboration with a tertiary eye hospital in India, CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously. CataractBot features multimodal support and multilingual capabilities. In an in-the-wild deployment study with 49 participants, CataractBot proved valuable, providing anytime accessibility, saving time, and accommodating diverse literacy levels. Trust was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
@@ 1. Introduction
The evolving landscape of healthcare is witnessing a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007). It signifies a departure from the traditional passive patient role toward a more informed, empowered, and engaged approach to healthcare, resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their health conditions, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery or radiotherapy, this demand for information becomes more crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Conversely, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
@@
Despite the abundance of available information sources, the digital age often overwhelms individuals with an excess of information, much of which is inaccurate or unreliable (Clarke et al., 2016a; Kinnersley et al., 2007). This poses a challenge in discerning trustworthy sources from misleading ones, leading to potential misconceptions. Additionally, patients frequently encounter difficulties finding information online at an appropriate level–ranging from oversimplified to overly technical–and harbor skepticism toward its reliability (Xiao et al., 2023). As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although locating them can be challenging (Civan-Hartzler et al., 2010). Therefore, patients primarily trust their doctor and the hospital staff responsible for their treatment, turning to them for any medical or logistical queries (Attfield et al., 2006; Clarke et al., 2016a). Studies examining patients’ information needs reveal a strong desire for doctor-endorsed health information (Tang and Newcomb, 1998). Study participants noted that being better informed increased their understanding of treatment plans, motivation to adhere to them, and satisfaction with hospital visits.
@@
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Even providing all potentially relevant information can pose challenges as it can overwhelm patients with cognitive overload (Christalle et al., 2019). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, subsequently alleviating the burden on doctors and medical staff.
To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
@@
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with the medical team at a tertiary eye hospital in Bangalore, India, we exemplified our approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized RAG (Retrieval-Augmented Generation) using a knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and updates to the knowledge base based on expert-provided edits to minimize future expert intervention. In particular, for high-stakes medical settings, access to reliable and trustworthy information is crucial, hence, each response that patients (or their attendants) receive undergoes scrutiny and potential editing by experts (doctors or patient coordinators). CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantaneously and subsequently providing verification at the experts’ convenience.
With that, the paper aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and their attendants) undergoing surgical treatment? (RQ2) How does the inclusion of an expert-in-the-loop influence the perceived trustworthiness of the information provided by the chatbot? (RQ3) How are such chatbots integrated into the workflow of doctors, patient coordinators, patients, and their attendants?
@@
To answer them, we conducted an in-the-wild deployment study involving 49 information seekers (19 patients, 30 attendants) using CataractBot for at least two weeks—a week before and after the surgery. Additionally, four doctors and two patient coordinators served as experts in this study. Our findings revealed patients and attendants highly appreciated CataractBot for its anytime accessibility, alleviating hesitations associated with power differences in asking questions, saving time by minimizing hospital visits/calls, and accommodating individuals with low literacy and tech proficiency through multilingual and multimodal support. Trust in the system was established through rigorous verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous knowledge base updates. Our main contributions are: (1) The conception and development of a specialized chatbot111Our code is available at https://github.com/microsoft/BYOeB., designed in close collaboration with doctors and hospital staff, to assist patients undergoing Cataract surgery with their information needs. (2) The deployment of this chatbot CataractBot in a real-world setting, among 31 Cataract surgery patients and their attendants at a tertiary eye hospital in India. Four doctors and two patient coordinators served as experts, actively verifying and refining responses generated by the chatbot. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, our study not only presents pivotal insights specific to CataractBot but also offers broader implications of LLM-powered expert-in-the-loop chatbots.
@@ 2. Related Work
Our work is primarily informed by two areas of relevant research: solutions to address patients’ information needs and conversational agents used in healthcare settings.
2.1. Addressing Patients’ Information Needs
Prior work has focused on understanding the diverse informational requirements of patients, particularly in the context of healthcare decision-making (Tseng et al., 2023; Xiao et al., 2023; Beisecker and Beisecker, 1990). Patients and their attendants exhibit a pattern of seeking information throughout the various stages of a patient’s journey, such as to prepare for doctor consultations (Radionova et al., 2023; Leong et al., 2018), to verify diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009), to assess the need for clinical interventions (Yang et al., 2023; Tseng et al., 2023), and to aid in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018). Studies emphasize that satisfying these diverse information needs enhances patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998).
To fulfill these diverse informational needs, patients, while trusting physicians for their clinical expertise, often turn to the Internet for the ease of access to information (Xiao et al., 2023). However, the challenge lies in identifying relevant and appropriate information, with instances of individuals consuming inaccurate information (Clarke et al., 2016b). Information seekers also resort to consulting experienced patients within their social network who can share valuable knowledge and insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010).
@@ 2.1. Addressing Patients’ Information Needs
Recent work in the field of Computer-Human Interaction has explored novel avenues to address the medical information gaps. For instance, Leong et al. (2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, while Wilcox et al. (2010) and Bickmore et al. (2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details based on their records. Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005). Patient-facing EHR services have received mixed reviews, with healthcare professionals expressing concerns about adding strain to already understaffed healthcare systems, while patients generally welcome the idea (Cajander and Grünloh, 2019b).
@@
In summary, patients require reliable, relevant, and timely information. However, challenges arise in accessing such information from busy or hesitant doctors and hospital staff. This dichotomy results in patients either receiving generic, unverified information from the internet and social circles or obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts. Our solution aims to strike a balance by providing a free, generic, high-quality answer instantaneously using Large Language Models (LLMs), which is later verified (and corrected if needed) by a medical professional with minimal increase in their workload.
@@ 2.2. Chatbots in Healthcare
Recently, researchers have been exploring chatbots as a means to provide easy and direct access to information, due to their ubiquity, minimal learning curve, and ability to offer a more personalized experience through natural language conversation (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020). Chatbots have been found to enhance the user experience by enabling users to engage in longer conversations and build complex queries message by message while retaining conversation context (Xiao et al., 2023). Stamenova et al. (2020) reveal that patients prefer messaging over video consultations, emphasizing that chatbots can help promote continuity in the treatment. Chatbots have been applied in various healthcare settings by the research community, including mental health support (Kumar et al., 2023; Jo et al., 2023), appointment scheduling, seeking information for chronic disease management (Montagna et al., 2023), assistance for new mothers (Yadav et al., 2019), and public health interventions to support socially isolated individuals (Jo et al., 2023). Commercial offerings, such as Babylon Health111Babylon Health: https://www.emed.com/uk, Ada222Ada: https://ada.com/, and Florence333Florence: https://florence.chat/, allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs, and the involvement of experts is limited to the data curation step (Xiao et al., 2023). Hence, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
@@
Beyond standalone chatbots, prior work has explored chat as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both the patient and the doctor to be simultaneously active, resembling an in-person consultation, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp, WeChat (Wang et al., 2020), or email enable communication between the patient and doctor based on their convenience, providing flexibility. Previous research indicates that both doctors and patients prefer asynchronous texts over instant calls (audio or video) (Stamenova et al., 2020; Johansson et al., 2020a).
In our work, we combine synchronous and asynchronous modes of communication, leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode. This hybrid approach represents a novel and effective method for addressing information needs in the healthcare domain.

### v2

@@ Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we develop CataractBot, an experts-in-the-loop chatbot powered by LLMs, in collaboration with an eye hospital in India. CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 55 participants, CataractBot proved valuable, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
@@ 1. Introduction
The evolving landscape of healthcare witnesses a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007), resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their symptoms, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery or radiotherapy, satisfying the demand for information becomes crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Similarly, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
@@
Despite the abundance of available information sources, the digital age often overwhelms individuals with an excess of information, much of which is inaccurate or unreliable (Clarke et al., 2016a; Kinnersley et al., 2007). This poses a challenge in discerning trustworthy sources from misleading ones. Patients also frequently encounter difficulties finding information online at an appropriate level, ranging from oversimplified to overly technical (Xiao et al., 2023). As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although identifying them can be challenging (Civan-Hartzler et al., 2010). Therefore, patients primarily trust their doctor and the hospital staff responsible for their treatment, turning to them for any medical or logistical queries (Attfield et al., 2006; Clarke et al., 2016a). Studies examining patients’ information needs reveal a strong desire for doctor-endorsed health information (Tang and Newcomb, 1998).
@@
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records (EHR) often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. However, providing all potentially relevant information risks overwhelming patients with cognitive overload (Christalle et al., 2019). Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, subsequently alleviating the burden on doctors and medical staff. To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
@@
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified our experts-in-the-loop chatbot approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (RAG) over a Knowledge Base (KB) curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and a KB update framework based on expert-provided edits to minimize future expert intervention. CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantly, and subsequently providing verification at the experts’ convenience.
Our work aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving 49 information seekers111We use the terms ‘participants’ or ‘information seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators. (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings revealed that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous KB updates.
@@
Our main contributions are: (1) The design and development of a novel experts-in-the-loop chatbot222Our code is publicly available at https://removed_for_anonymity., designed in collaboration with doctors and hospital staff, to assist patients undergoing cataract surgery with their information needs. (2) The deployment of our bot in a real-world setting, among 49 cataract surgery patients/attendants at a tertiary eye hospital in India. Four doctors and two patient coordinators served as experts, actively verifying and correcting responses generated by the bot. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
@@ 2. Related Work
Our work is primarily informed by two areas of relevant research: solutions addressing patients’ information needs and the use of conversational agents, including those powered by LLMs, in healthcare settings.
2.1. Patients’ Information Needs
Patients and their attendants seek information throughout the various stages of their medical journey, such as preparing for doctor consultations (Radionova et al., 2023; Leong et al., 2018), verifying diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009; Christalle et al., 2019), assessing the need for clinical interventions (Yang et al., 2024; Tseng et al., 2023), and aiding in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018). Studies emphasize that satisfying these diverse information needs improves patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998). Patients primarily trust doctors with their medical queries (Ding et al., 2020). However, traditional in-person consultations pose difficulties, including high costs (Bodenheimer and Smith, 2013), long wait times, and a strong hierarchy in the patient-physician relationship (Fochsen et al., 2006). Patients often miss opportunities to ask additional (clarification) questions during these brief encounters, as they face information overload (Ferré et al., 2020; Kinnersley et al., 2007). Doctors also struggle with time due to staff shortages straining healthcare ecosystems (Bodenheimer and Smith, 2013; Akerkar and Bichile, 2004). Thus, patients and caregivers are increasingly turning to the Internet for easier access to medical information (Xiao et al., 2023; Chandwani and Kulkarni, 2016). However, the challenge lies in identifying relevant and appropriate information. Individuals struggle to understand complex health information or may even consume inaccurate information, making the process frustrating (Zhang, 2010; Arora et al., 2008; Clarke et al., 2016b).
Information seekers may also consult experienced patients within their social networks who can share valuable insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010). In the absence of such known individuals, people may resort to social media, which is not only ineffective for quality health support due to widespread misinformation and poor management (Malki et al., 2024), but also risky in terms of personal privacy (De Choudhury et al., 2014). Chandwani and Kulkarni (2016) found that although most doctors regard the practice of seeking health information online as an affront to their authority, some see it as inevitable–and even as an opportunity to elevate patients’ health literacy through technology.
@@ 2.1. Patients’ Information Needs
Recent work in the field of Computer-Human Interaction has explored novel avenues to address these medical information gaps. For instance, Leong et al. (2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, Wilcox et al. (2010) and Bickmore et al. (2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details, and Pfeifer Vardoulakis et al. (2012) propose using mobile phones to present dynamic, interactive reports on patients’ progress and care plans throughout their emergency department stay. Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005; Chung, 2017). Patient-facing EHR services have received mixed reviews, with healthcare professionals expressing concerns about adding strain to already understaffed healthcare systems, while patients generally welcome the idea (Cajander and Grünloh, 2019b).
@@
In summary, patients need reliable, relevant, and timely information. However, accessing this from busy doctors and hospital staff is challenging. This dichotomy results in patients either receiving generic, unverified, and hard-to-understand information from the internet and social circles or if lucky, obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts. Our solution aims to strike a balance by providing a generic, high-quality answer instantly using LLMs, which is later verified by a medical professional with minimal increase to their workload.
@@ 2.2. Chatbots in Healthcare
Chatbots enable natural language conversation, resulting in a minimal learning curve and offering a personalized experience (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020; Chung, 2017; Desai and Chin, 2023). They enable users to build complex queries message by message while retaining conversation context (Xiao et al., 2023). Researchers have applied chatbots in various healthcare settings (Wutz et al., 2023), including mental health support (Kumar et al., 2023; Jo et al., 2023; Ma et al., 2023), appointment scheduling (Ittarat et al., 2023), seeking information for chronic disease management (Montagna et al., 2023), and assistance for new mothers (Yadav et al., 2019). Commercial chatbots, such as Babylon Health, Ada, and Florence333Babylon Health: https://www.emed.com/uk. Ada: https://ada.com/. Florence: https://florence.chat/., allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs. However, the involvement of doctors in these bots is limited to the data curation step (Xiao et al., 2023). Consequently, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
LLMs have been criticized for hallucinating and not displaying the reason behind their outputs (Fogel and Kvedar, 2018). Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (RAG) (Lewis et al., 2020), and few-shot learning. Prior work suggests that healthcare professionals are more likely to adopt AI tools when they perceive them as ‘partners’ that enhance their workflow and expertise (Henry et al., 2022). Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated chatbots could potentially combine the best of both worlds, a concept we propose in CataractBot.
Given our specific focus on cataract surgery, we also examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
@@
Beyond chatbots, prior work has explored “chat” as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both the patient and the doctor to be simultaneously active, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp (Yadav et al., 2019), WeChat (Wang et al., 2020), SMS (Perrier et al., 2015), email (Makarem and Antoun, 2016), or web portals (Mamykina et al., 2008; Tuli et al., 2018) enable communication at the convenience of both parties, providing flexibility. Previous research indicates that both doctors and patients prefer asynchronous texts over instant audio/video calls (Stamenova et al., 2020; Johansson et al., 2020a), despite lacking instant feedback (Li et al., 2023). In our work, we combine both approaches—leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode. This hybrid approach represents a novel and effective method for addressing information needs in the healthcare domain.
Finally, the use of AI in real-world healthcare deployments is limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and validated its utility to understand its integration into clinical workflows.
3. Formative Study
The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital. To identify key chatbot requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023. From these interviews, we aimed to understand their current workflow, the types of questions patients ask, and their concerns regarding patients’ knowledge gaps. Participants had an average age of 38.58.8 years, with 11.07.5 years of work experience. They participated in the study voluntarily without compensation. The interviews were conducted in English and lasted around 45 minutes each. All sessions were audio-recorded and transcribed by the interviewer. We conducted a thematic analysis to identify these key design requirements for the chatbot.
Accuracy: The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People will have weird questions… like ‘when can I do ayurvedic nasal therapy post-op?’.” Questions related to diet require awareness of Indian food options for accurate guidance. To ensure accurate and contextual responses, our chatbot relies on a Knowledge Base (KB) curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
Trustworthiness: Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or friends, they often confirm that information with their doctor. E.g., a doctor noted, “They come and ask us… my friends have told me this… is it correct or not?” To ensure trustworthiness in CataractBot we decided that (a) each response will be verified and, if necessary, edited by doctors and patient coordinators, and (b) patients (or their attendants) will be explicitly notified when this expert verification happens. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
Timeliness: Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to busy schedules of doctors and coordinators, immediate responses are not always feasible. This often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using a custom KB curated by hospital staff. Subsequently, these responses gets reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. To ensure swift verification of every query, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
Usability: After cataract surgery, patients are expected to reference their discharge summaries for any queries, but this can be challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role immediate family members play in monitoring the patient’s healthcare. Coupled with doctors’ input (“Hardly 10% of questions are asked by patients; attendants are the one who ask.”), this underscored the necessity for the bot to be usable by the four key stakeholders of the cataract surgery ecosystem: patients, their attendants, doctors, and patient coordinators. To cater to diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, as suggested by doctors and patient coordinators, we made the chatbot multilingual and multimodal similar to (Jain et al., 2018; Findlater et al., 2009). We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022), despite the richer feature set of Telegram or the development flexibility of a custom chatbot. This choice aims to minimize the learning curve and ensure quick onboarding without requiring additional app installations. Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal and ‘symbolic’ messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s KB, minimizing similar edits in future.